OpenStudy (anonymous):

Write the expression: 1. the sum of 6 and t = 6+ t PLEASE HELP STEP BY STEP

OpenStudy (anonymous):

I think you did it already.

OpenStudy (anonymous):

Write the expression: 1. the sum of 6 and t = 6+ t a. the product of s and 12

OpenStudy (anonymous):

6+t and s*12

OpenStudy (anonymous):

* what does that mean?

OpenStudy (anonymous):

sum of 6 and t just means 6+t and the product of 12 and s is just 12*s

OpenStudy (anonymous):

what is *

OpenStudy (anonymous):

what are you trying to find out

OpenStudy (anonymous):

Write the expression: 1. the sum of 6 and t = 6+ t a. the product of s and 12

OpenStudy (anonymous):

1. 6+t=n 2. s*12=n i used n as the variable filling in for the answer